OpenStudy (anonymous):

Mary's age is 2/3 that of Peter's.Two years ago Mary's age was 1/2 of what peter's age will be in 5 year's time.How old is Peter now?

OpenStudy (anonymous):

@sirm3d answer plz

OpenStudy (anonymous):

urgent help

OpenStudy (anonymous):

@KatieMO

OpenStudy (anonymous):

hmm... give me a min...

OpenStudy (anonymous):

Yeah, I have no idea... what subject is this??? lol

OpenStudy (anonymous):

peter age=x marie age=2/3x x+5=.5(2/3x-2) x-.33x=-5-1 .67x=-6

OpenStudy (anonymous):

P = 1.5 M M -2 = .5 ( P + 5 ) -2 M -2 = .5P + 2.5 - 2 M -2 = ,5 * 1.5M + .5 .5M = 3 M = 6 yr old =>P = 9 yr old